# Multi-Grained Query-Guided Set Prediction Network for Grounded Multimodal Named Entity Recognition

### v1

@@ Abstract.
Grounded Multimodal Named Entity Recognition (GMNER) is an emerging information extraction (IE) task, aiming to simultaneously extract entity spans, types, and entity-matched bounding box groundings in images from given sentence-image pairs data. Recent unified methods employing machine reading comprehension (MRC-based) frameworks or sequence generation-based models face challenges in understanding the relationships of multimodal entities. MRC-based frameworks, utilizing human-designed queries, struggle to model intra-entity connections. Meanwhile, sequence generation-based outputs excessively rely on inter-entity dependencies due to pre-defined decoding order. To tackle these, we propose a novel unified framework named Multi-grained Query-guided Set Prediction Network (MQSPN) to learn appropriate relationships at intra-entity and inter-entity levels. Specifically, MQSPN consists of a Multi-grained Query Set (MQS) and a Multimodal Set Prediction Network (MSP). MQS combines specific type-grained and learnable entity-grained queries to adaptively strengthen intra-entity connections by explicitly aligning visual regions with textual spans. Based on solid intra-entity modeling, MSP reformulates GMNER as a set prediction, enabling the parallel prediction of multimodal entities in a non-autoregressive manner, eliminating redundant dependencies from preceding sequences, and guiding models to establish appropriate inter-entity relationships from a global matching perspective. Additionally, to boost better alignment of two-level relationships, we also incorporate a Query-guided Fusion Net (QFNet) to work as a glue network between MQS and MSP. Extensive experiments demonstrate that our approach achieves state-of-the-art performances in widely used benchmarks. Notably, our method improves 2.83% F1 in the difficult fine-grained GMNER benchmark.  [^1] Code is available at https://github.com/tangjielong928/mqspn.
Grounded Multimodal Named Entity Recognition, Multimodal Named Entity Recognition, Set Prediction, Unified Model
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXprice: 15.00ccs: Computing methodologies Information extraction
1. Introduction
Extracting multimodal entity information, including entity spans, types, and entity-matched bounding box groundings through unified models is an emerging solution of Grounded Multimodal Named Entity Recognition (GMNER) (Yu et al., 2023). Although recent unified works have made progress by employing machine reading comprehension (MRC-based) frameworks and sequence generation-based models, limitations still exist. The MRC-based frameworks (Jia et al., 2022, 2023) utilize human-designed queries to provide untrainable prior instructions, struggling to model the connections between textual spans and visual regions, and leading to misdistinguishing same-type entity regions (‘Same-Type Error’ in Figure 1 (b)). On the other hand, Sequence generation-based methods (Yu et al., 2023; Wang et al., 2023) autoregressively decode all span-type-region triples one by one in predefined sequence order resulting in current outputs highly sensitive to errors in preceding sequences (‘Sequence-Sensitive Error’ in Figure 1 (c)).
Figure 1. The comparison of existing approaches and our MQSPN. (a) The task definition of GMNER includes inputs of image-text pairs and outputs of multiple span-type-region triplets. (b) The MRC-based approach employs human-designed queries to instruct the model in joint learning, leading to ‘Same-Type Error’ due to intra-entity misunderstanding (c) The sequence generation-based approach decodes span-type-region triples one by one, leading to ‘Sequence-Sensitive Error’ due to inter-entity overreliance. (d) Our MQSPN model appropriate two-level relationships with learnable query set and set prediction.
We argue that it is difficult for these unified works to understand two-level relationships of multimodal entities. The ‘Same-Type Error’ in the MRC-based framework presents a mismatch between spans and regions within a single multimodal entity and will not influence other entities which we call intra-entity relationship misunderstanding. Similarly, the ‘Sequence-Sensitive Error’ reveals that sequence generation-based models overly rely on dependencies between different multimodal entities which we define as inter-entity relationship overreliance. To address these issues, we propose a novel unified framework Multi-grained Query-guided Set Prediction Network (MQSPN) to focus on modeling appropriate relationships at intra-entity and inter-entity levels.
MQSPN mainly consists of Multi-grained Query Set (MQS), Multimodal Set Prediction Network (MSP), and Query-guided Fusion Net (QFNet). The main motivation of MQS is to adaptively learn intra-entity relationships. Instead of human-designed queries in the MRC-based framework, MQS adopts a set of learnable queries (Li et al., 2023b; Gong et al., 2024) (denoted as entity-grained query) for aligning visual regions with textual span. However, simply learnable queries are insufficient to detect regions and spans due to the lack of initialization of prior semantics. Inspired by Mask Language Modeling (Devlin et al., 2019), where the prediction of the [MASK] token represents the corresponding semantic embedding in natural language, we feed a prompt with masked type statement into vanilla BERT to build type-grained queries with rich type-specific initial semantics. We then construct MQS by associating a learnable entity-grained query with a type-grained query to a multi-grained query pair. Notably, unlike input single query during each iteration in the MRC-based framework, multiple query pairs are jointly employed in MQS to learn entity-grained query semantics for different multimodal entities and establish explicit intra-entity connections.
@@ 1. Introduction
Based on solid intra-entity modeling, we further apply MSP to explore suitable inter-entity relationships. Different from previous sequence generation-based methods, MSP reformulates GMNER as set predictions (Tan et al., 2021; Shen et al., 2022), as shown in the grey box of Figure 1(d). With the input of multiple queries from MQS, MSP parallelly predicts a set of multimodal entities in the non-autoregressive manner without the need for a preceding sequence. Given the set of predicted entities and ground truths, the objective of MSP is to find the optimal bipartite matching between them with minimal global matching cost. Additionally, since the number of queries is greater than the total quantity of gold entities, it is not easy to evaluate the predicted entities according to the gold entities. To address this, we first copy the ground truths to the size of MQS and then design a loss function based on bipartite matching. Finally, optimal matching can be efficiently solved by the off-the-shelf Hungarian Algorithm (Kuhn, 1955). In this manner, the inference of MSP will not depend on redundant dependencies dictated by a predefined decoding order, thereby guiding models to establish suitable inter-entity relationships from a global matching perspective.
@@
Besides, since it has been validated (Chen et al., 2021) that directly fusing textual and visual information will impair MNER model performance due to the introduction of noise from irrelevant visual regions, we also propose a QFNet between MQS and MSP to filter this noisy information, thereby boosting better alignment of two-level relationships. Unlike direct fusion methods (Yu et al., 2020; Zhang et al., 2021b; Wu et al., 2023), QFNet employs queries as intermediaries to facilitate the separate integration of textual and visual region representations. It incorporates three interaction mechanisms: Query-text Cross-Attention (QCT), Query-region Prefix Integration (QPI), and Similarity-aware Aggregator (SAG). QCT and QPI seamlessly integrate information across different modalities, while SAG mitigates the noise caused by misaligned regions and textual spans. More details are illustrated in Section 3.5. Our contributions could be summarized as follows:
@@
  • We rethink the essence of existing unified GMNER methods’ weaknesses from a new perspective, two-level relationships (intra-entity and inter-entity), and propose a unified framework MQSPN to solve intra-entity relationship misunderstanding and inter-entity overreliance.
@@
  • Extensive experiments on two Twitter benchmarks illustrate that our method outperforms existing state-of-the-art (SOTA) methods, especially in the difficult fine-grained GMNER benchmark. The ablation study also validates the effectiveness of each designed module.
2. Related Work
2.1. Multimodal NER and Grounded MNER
Exiting multimodal named entity recognition (MNER) models primarily focused on how to utilize visual information to assist textual models in entity extraction, which attempted to design reasonable cross-modal fusion mechanisms (Yu et al., 2020; Zhang et al., 2021b; Wu et al., 2023; Chen et al., 2022a), explore vision-language alignment(Xu et al., 2022; Wang et al., 2022b; Chen et al., 2023) or leverage external knowledge-augmented methods (Wang et al., 2022a, c; Li et al., 2023a) to enhance the model’s generalization capabilities. To facilitate the construction of a multimodal knowledge graph, a more challenging task named Grounded MNER (GMNER) is proposed to additionally output the bounding box coordinates of named entities within the image. To further expansion, Wang et al. (Wang et al., 2023) introduced fine-grained entity types in the GMNER task. Yu et al. (Yu et al., 2023) combined the previously superior MNER models with an off-the-shelf object detection toolkit to propose a series of strong baselines in the pipeline manner. Li et al (Li et al., 2024). and Ok et al (Ok et al., 2024). respectively utilize large language models (LLMs) and Wikipedia retrieval to acquire additional knowledge related to entities, thereby improving the pipeline framework performance. To tackle the error propagation issue, H-Index (Wang et al., 2023) and TIGER (Yu et al., 2023) formulate GMNER as a sequence generation task and use the BART (Yan et al., 2021) or T5 (Raffel et al., 2020) models to generate the span-type-region triples. MNER-QG (Jia et al., 2023) formulate GMNER as a machine reading comprehension task and utilize manually constructed type-specific queries to concurrently guide the model in performing both query grounding and entity span extraction. However, sequence generation-based methods autoregressively decode span-type-region triples in predefined sequence order, which excessively rely on inter-entity dependencies and influence the model’s generalization capability. As for the MRC-based methods, the manually constructed queries are semantically impoverished and struggle to model robust intra-entity connections. In this work, we model the GMNER task as set prediction in a non-autoregressive manner, which maintains suitable inter-entity relationships in a global matching perspective. Meanwhile, we replace the manually constructed queries with a learnable multi-grained query set, in which fine-grained entity semantics and intra-entity connection can be learned.
2.2. Set Prediction
@@
Set prediction is a well-known machine learning technique where the goal is to predict an unordered set of elements. It is widely used in various applications such as object detection (Carion et al., 2020), point cloud classification (Qi et al., 2017), and multi-label classification (Yang et al., 2019). Unlike traditional prediction tasks that output a single value or a predetermined sequence, set prediction models generate a collection of predictions without any inherent order. In the field of named entity recognition (NER), previous state-of-the-art methods (Yan et al., 2021) utilize autoregressive sequence generation models to output entities in a predefined sequence order, which violates the permutation invariance of predicted entities. To tackle this issue, Tan et al. (Tan et al., 2021) reformulate the nested NER task as set prediction and propose an innovative sequence-to-set model. Subsequently, Shen et al. (Shen et al., 2022) associate each entity with a learnable instance query to capture the location and type semantics, training with a dynamic label assignment mechanism. Recently, Chen et al. (Chen et al., 2024) extended set prediction to other sequence labeling tasks such as slot filling and part-of-speech (POS) Tagging to enhance performance and inference speed. Different from these methods, we introduce the MQSPN, which exploit the set prediction paradigm in a new GMNER area to model the two-level relationship of multimodal entities.
2.3. Visual Grounding
@@
Visual Grounding (VG) aims to detect the most relevant visual region based on a natural language query, i.e., phrase, category, or sentence. Most existing works can be divided into two branches. The first branch utilizes the one-stage detector, such as DETR (Carion et al., 2020) and YOLO (Redmon and Farhadi, 2018), to directly output object bounding boxes in an end-to-end manner. The second branch first generates candidate regions with some region proposal methods, such as Region Proposal Network (RPN) (Girshick, 2015) and selective search (Uijlings et al., 2013), and then selects the best-matching region based on language query. There are two main distinctions between VG and EG in the GMNER task. (1) VG models are typically trained on coarse-grained datasets, while EG is based on fine-grained named entities. (2) The language queries in the training set of VG can be grounded to corresponding visual regions, while most of the named entities in EG are ungroundable. These discrepancies make it difficult to train a solid one-stage visual detector. Hence, our work follows the two-stage paradigm and exploits VinVL (Zhang et al., 2021a) as a class-agnostic RPN.
Figure 2. (a). Overview of our MQSPN. The construction of Multi-grained Query Set (MQS) consists of a Prompt-based Type-grained Query Generator (PTQG) and a Learnable Entity-grained Query (LEQ) Set. The Multimodal Set Prediction Network (MSP) contains bipartite matching with three prediction heads, including Candidate Regions Matching (CRM), Span Boundary Localization (SBL), and Entity Classification(EC). The Query-guided Fusion Net (QFNet) is shown on the right (b). SAG denotes the Similarity-aware Aggregator. FGF is the fine-grained fusion module.
3. Our Method
3.1. Task Formulation
In this work, we approach GMNER as a set prediction, where the multimodal entities with their corresponding regions are represented as a set of quadruples. Specifically, given a training sample $(X,I,Y)$, where $X=(w_{1},w_{2},...,w_{n})$ is the input sequence of length $n$, $I$ is the corresponding image, and $Y$ is the set of gold multimodal entities:
(1) $Y=\{\left(Y_{1}^{s},Y_{1}^{e},Y_{1}^{t},Y_{1}^{r}\right),\ldots,\left(Y_{m}^{s},Y_{m}^{e},Y_{m}^{t},Y_{m}^{r}\right)\},$
@@ 3.1. Task Formulation
where $(Y_{i}^{s},Y_{i}^{e},Y_{i}^{t},Y_{i}^{r})$ denote the $i$-th quadruple, $Y_{i}^{s}\in[0,n-1]$ and $Y_{i}^{e}\in[0,n-1]$ are the start and end boundary indices of the $i$-th target entity span. $Y_{i}^{t}$ refers to its corresponding entity type, and $Y_{i}^{r}$ denotes visually grounded region of the $i$-th target entity. Note that if the target entity cannot be grounded in the given image $I$, $Y_{i}^{r}$ is $None$; otherwise, $Y_{i}^{r}$ consists of a 4-D spatial feature including the top-left and bottom-right positions of the grounded bounding box, i.e.,$(Y_{i}^{r^{x_{1}}},Y_{i}^{r^{y_{1}}},Y_{i}^{r^{x_{2}}},Y_{i}^{r^{y_{2}}})$. To enable the set prediction, we assign $u(u>m)$ learnable queries set for each training sample, which can be denoted as $Q=\mathbb{R}^{u\times h}$. Each query $q\in Q$ (a vector of size $h$) is responsible for extracting one corresponding multimodal entity quadruple $(Y_{i}^{s},Y_{i}^{e},Y_{i}^{t},Y_{i}^{r})$. Consequently, we can define the GMNER task as follows: given an input sentence $X$ and corresponding image $I$, the objective is to match the predicted quadruples based on learnable queries set $Q$ with the gold quadruples set ${Y}$.
3.2. Overview
As illustrated in Figure 2, we present a set prediction-based method named MQSPN with four different components, consisting of the Feature Extraction Module, Multi-grained Query Set (MQS), Query-guided Fusion Net (QFNet) and Multimodal Set Prediction Network (MSP). During training, the Feature Extraction Module first generates candidate region proposals and then encodes them with the given text to obtain textual and visual features. Subsequently, learnable MQS, textual features, and visual features pass through QFNet to perform modality fusion and filter noisy information. Finally, the last hidden states of three features are fed into MSP to perform joint training via optimal bipartite matching. Notably, MQS guides the model to learn explicit intra-entity connections, while MSP formulates inter-entity relationships from a global matching perspective. Details of the model are provided as follows.
3.3. Feature Extraction Module
@@
Text Representation. Given the input sentence $X$, textual encoder BERT (Devlin et al., 2019) is used to tokenize it into a sequence of word embeddings $H_{T}=([CLS],e_{1},...,e_{n},[SEP])$, where $e_{i}\in\mathbb{R}^{h},$ $h$ is the hidden dimension, $[CLS]$ and $[SEP]$ are special tokens of the beginning and end positions in word embeddings.
Visual Representation. Given the input image $I$, we utilize the class-agnostic region proposal network(RPN) VinVL (Zhang et al., 2021a) to obtain all candidate regions. Following the work of (Yu et al., 2023) and (Wang et al., 2023), we also retain the top-$k$ region proposals as our candidate regions, using the last hidden layer embeddings of VinVL as the initial visual representation for the candidate regions, denoted as $\mathbf{R}=\{\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{k}\},$ where $\mathbf{r}_{i}\in\mathbb{R}^{2048}.$ is the initial visual feature of the $i$-th candidate region. A linear projection layer is then used to map the initial regional embeddings to dimensions consistent with the textual representations, and the candidate regional representations are denoted as $\mathbf{V}=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\},$ where $\mathbf{V}_{i}\in\mathbb{R}^{h}.$
@@
To match those entities that are ungroundable in the images, we construct a special visual token embedding $\mathbf{v}_{[ug]}\in\mathbb{R}^{h}$ by feeding a blank image $\mathcal{T}$ into ViT-B/32 (Dosovitskiy et al., 2020) from pre-training CLIP (Radford et al., 2021). Finally, $\mathbf{v}_{[ug]}$ and $\mathbf{V}$ are concatenated to serve as the final visual representation $H_{V}\in\mathbb{R}^{(k+1)\times h}$, where $k$ is the number of candidate regions.
3.4. Multi-grained Query Set Construction
@@
Previous MRC-based methods attempted to incorporate prior knowledge using manually constructed query statements. However, these fixed and coarse-grained queries struggled to provide distinguishable features for same-type entities, hindering the modeling of intra-entity connections. In this section, we propose a learnable multi-grained query set to overcome this issue.
@@
(2) $H_{Q}^{o}=BERT\left(Prompt\right)_{[\text{MASK}]}$
@@
Learnable Entity-grained Query Set. Entity-grained queries are randomly initialized as learnable input embeddings $H_{Q}^{e}\in\mathbb{R}^{u\times h}$. During training, the entity-grained semantics and corresponding relationships between candidate regions and entity spans can be learned automatically by these embeddings. To ensure that type-grained query embedding and entity-grained query embedding have the same dimensions, we replicated the former $d$ times (note that $p<u$ and $d=u/p$). Then the multi-grained query embedding is given by:
(3) $H_{Q}\mathbf{}=H_{Q}^{e}\oplus\left[H_{Q}^{o}\right]^{d}$
@@
3.5. Query-guided Fusion Net
Previous multimodal fusion approaches mainly suffered from semantic discrepancies and heterogeneity issues across different modalities. The direct fusion of textual and visual information could impair model performance. Different from previous fusion methods, we use queries as intermediaries to guide the integration of textual representations and visual region representations respectively, thereby reducing the introduction of noise from irrelevant information. To be specific, we present two interaction mechanisms (query-text cross-attention interaction and query-region prefix integration) to seamlessly integrate information across different modalities. To achieve fine-grained alignment between queries and both textual tokens and visual regions, we propose a similarity-aware aggregator that reduces noise introduced by irrelevant candidate regions. We will provide detailed explanations of each module in the following section.
@@ 3.5. Query-guided Fusion Net
Query-text Cross-attention Interaction. As shown in Figure 2 (b), the query set $H_{Q}$ (obtained from Section 3.4 and each query treated as a special textual token) and the textual sequence $H_{T}$ (obtained from Section 3.3) are fed into the transformer-based architecture (Vaswani et al., 2017). The cross-attention mechanism is used to fuse these unimodal features, which can be expressed as follows:
(4) $\mathrm{C}T_{Q\to T}(H_{Q},H_{T})=\mathrm{softmax}\left(\frac{H_{Q}W_{Q_{q}}W_{T_{k}}^{\top}H_{T}^{\top}}{\sqrt{d_{k}}}\right)H_{T}W_{T_{v}}$
$\mathrm{C}T_{T\to Q}(H_{T},H_{Q})=\mathrm{softmax}\left(\frac{H_{T}W_{T_{q}}W_{Q_{k}}^{\top}H_{Q}^{\top}}{\sqrt{d_{k}}}\right)H_{Q}W_{Q_{v}}$
where $\mathrm{C}T_{{\alpha\to\beta}}(\cdot)$ is the cross-attention calculation from $\alpha$ to $\beta$. $W_{\alpha_{q}}\in\mathbb{R}^{d_{\alpha}\times d_{k}}$, $W_{\beta_{k}}\in\mathbb{R}^{d_{\beta}\times d_{k}}$, and $W_{\beta_{v}}\in\mathbb{R}^{d_{\beta}\times d_{v}}$ represent the matrices used to projecting queries, keys, and values respectively.
Query-region Prefix Integration. Inspired by prefix tuning (Li and Liang, 2021) and its successful applications (Chen et al., 2022a, b) in multimodal fusion, we propose query-region prefix integration to reduce the semantic bias and heterogeneity between multimodal data. Specifically, we use the query set features as a visual prefix, inserting them into the candidate region features at keys and values layers of each multi-head attention in the vision transformer (Dosovitskiy et al., 2020). First, the $(l-1)$-th layer candidate region features $H_{V}^{l-1}$ is projected into the query, key, and value vectors :
(5) $q_{V}^{(l)}=H_{V}^{(l-1)}W_{V_{q}}^{(l)},k_{V}^{\left(l\right)}=H_{V}^{\left(l-1\right)}W_{V_{k}}^{\left(l\right)},v_{V}^{\left(l\right)}=H_{V}^{\left(l-1\right)}W_{V_{v}}^{\left(l\right)}$
(6) $\{\rho_{k}^{l},\rho_{v}^{l}\}=H_{Q}^{l-1}W_{\rho}^{l},$
(7) $PI^{\left(l\right)}=\mathrm{softmax}(\frac{q_{V}^{\left(l\right)}[\rho_{k}^{\left(l\right)};k_{V}^{\left(l\right)}]^{T}}{\sqrt{h}})[\rho_{v}^{\left(l\right)};v_{V}^{\left(l\right)}]$

### v2

@@ Abstract
Grounded Multimodal Named Entity Recognition (GMNER) is an emerging information extraction (IE) task, aiming to simultaneously extract entity spans, types, and corresponding visual regions of entities from given sentence-image pairs data. Recent unified methods employing machine reading comprehension or sequence generation-based frameworks show limitations in this difficult task. The former, utilizing human-designed queries, struggles to differentiate ambiguous entities, such as Jordan (Person) and off-White x Jordan (Shoes). The latter, following the one-by-one decoding order, suffers from exposure bias issues. We maintain that these works misunderstand the relationships of multimodal entities. To tackle these, we propose a novel unified framework named Multi-grained Query-guided Set Prediction Network (MQSPN) to learn appropriate relationships at intra-entity and inter-entity levels. Specifically, MQSPN consists of a Multi-grained Query Set (MQS) and a Multimodal Set Prediction Network (MSP). MQS explicitly aligns entity regions with entity spans by employing a set of learnable queries to strengthen intra-entity connections. Based on distinct intra-entity modeling, MSP reformulates GMNER as a set prediction, guiding models to establish appropriate inter-entity relationships from a global matching perspective. Additionally, we incorporate a query-guided Fusion Net (QFNet) to work as a glue network between MQS and MSP. Extensive experiments demonstrate that our approach achieves state-of-the-art performances in widely used benchmarks [^1] Code is available at https://github.com/tangjielong928/mqspn..
Introduction
To effectively comprehend and manage vast amounts of multimodal content from social media, recent research (Yu et al. 2023) proposes a nascent multimodal information extraction (IE) task named Grounded Multimodal Named Entity Recognition (GMNER), which aims to extract multimodal entity information, including entity spans, types, and corresponding regions of entities, from image-text pairs. Prior studies decompose the GMNER task into several subtasks such as Multimodal Named Entity Recognition (Zhang et al. 2021a) and Visual Grounding (Redmon and Farhadi 2018), adopting the pipeline approach to solve it, which leads to serious error propagation. To address this problem, recent research paradigm has transferred to detect span-type-region triplets with unified model by formulating GMNER as machine reading comprehension (MRC) (Jia et al. 2023) or sequence generation (Yu et al. 2023; Wang et al. 2023).
Despite their remarkable performance, limitations still exist. The MRC-based frameworks utilize human-designed queries as prior instructions to simultaneously guide entity recognition and entity grounding, struggling to distinguish different ambiguous entities. For example in Figure 1 (a), with the input of multiple fixed person queries like “Please extract person: People’s name…”, the model incorrectly detects off-White x Jordan (Shoes) as Jordan (Person) and assigns the wrong region originally belonging to Kevin Durant. On the other hand, sequence generation-based methods suffer from exposure bias. They autoregressively decode span-type-region triples one by one in predefined sequence order, resulting in the prediction of off-White x Jordan region highly sensitive to errors in preceding Kevin Durant detection in Figure 1 (b).
Figure 1: The comparison of existing approaches and our MQSPN. (a) The MRC-based methods trap in entity-ambiguous issue due to intra-entity misunderstanding. (b) The sequence generation-based methods suffer from exposure bias issue due to inter-entity overreliance. (c) Our MQSPN model appropriate two-level relationships with learnable query set and set prediction.
In our perspective, the essence of these errors is the inappropriate modeling of multimodal entity relationships. (1) MRC-based frameworks overlook the distinctions of intra-entity relationships due to their reliance on duplicate and untrainable queries that cannot distinguish intrinsic features of ambiguous entities. (2) Sequence generation-based methods excessively rely on inter-entity relationships between different multimodal entities, where the current output is vulnerable to previous predictions. Therefore, we propose a novel unified framework named MQSPN, mainly consisting of learnable Multi-grained Query Set (MQS) and Multimodal Set Prediction Network (MSP) to focus on modeling appropriate relationships at intra-entity and inter-entity levels.
We maintain that modeling intra-entity connections is foundational; only when a model is capable of distinguishing individual entities can it further model inter-entity relationships effectively. Hence, we propose MQS to adaptively learn intra-entity connections. Instead of human-designed queries in the MRC-based framework, MQS adopts a set of learnable queries (Li et al. 2023; Gong et al. 2024) (denoted as entity-grained query) to perform joint learning in span localization, region grounding, and entity classification for different entities, which enforces queries to learn distinguishable features and automatically establish explicit intra-entity connections. However, simply learnable queries are insufficient to detect regions and spans due to the lack of semantics information. To tackle this, we feed a prompt with masked type statement into vanilla BERT to build type-grained queries with type-specific semantics. Finally, each multi-grained query is constructed by integrating a learnable entity-grained query with a type-grained query.
@@ Introduction
Based on distinct intra-entity modeling, we further apply MSP to explore suitable inter-entity relationships. Different from previous sequence generation-based methods, MSP reformulates GMNER as set predictions (Tan et al. 2021; Shen et al. 2022). As shown in Figure 1 (c), with one-time input of learnable MQS, MSP parallelly predicts a set of multimodal entities in the non-autoregressive manner without the need for a preceding sequence. The training objective of MSP is to find the optimal bipartite matching with minimal global matching cost, which can be efficiently solved by the off-the-shelf Hungarian Algorithm (Kuhn 1955). In this manner, the inference of MSP will not depend on redundant dependencies dictated by a predefined decoding order, thereby guiding models to establish suitable inter-entity relationships from a global matching perspective.
@@
Besides, since directly fusing textual features with irrelevant visual features will impair model performance (Chen et al. 2021), we further propose a QFNet between MQS and MSP to filter this noisy information, thereby boosting better alignment of two-level relationships. Unlike direct fusion methods (Yu et al. 2020; Zhang et al. 2021b; Wu et al. 2023), QFNet employs queries as intermediaries to facilitate the separate integration of textual and visual region representations. More details are illustrated in the following section. Our contributions could be summarized as follows:
@@
  • We delve into the essence of existing unified GMNER methods’ weaknesses from a new perspective, two-level relationships (intra-entity and inter-entity), and propose a unified framework MQSPN to adaptively learn intra-entity relationships and establish inter-entity relationships from global optimal matching view.
@@
  • Extensive experiments on two Twitter benchmarks illustrate that our method outperforms existing state-of-the-art (SOTA) methods. The ablation study also validates the effectiveness of each designed module.
Related Work
Grounded Multimodal Named Entity Recognition. Previous multimodal named entity recognition (MNER) models (Yu et al. 2020; Zhang et al. 2021b; Chen et al. 2022a; Zhang et al. 2021a; Wang et al. 2022b) primarily focused on how to utilize visual information to assist textual models in entity extraction. Grounded MNER (GMNER) (Yu et al. 2023) is proposed to additionally output the bounding box coordinates of named entities within the image. The taxonomy of previous GMNER works encompasses two branches, namely pipeline manner and unified manner. Pipeline methods (Li et al. 2024a; Ok et al. 2024) decompose the GMNER task into several subtasks like MNER, entity entailment and visual grounding (Wang et al. 2022a). To tackle the error propagation issue, unified methods formulate GMNER as an end-to-end machine reading comprehension (Jia et al. 2023, 2022) or sequence generation task (Wang et al. 2023; Yu et al. 2023; Li et al. 2024b). Different from them, our MQSPN reformulate GMNER as set prediction to learn appropriate intra-entity and inter-entity relationships.
@@
Set Prediction. Set prediction is a well-known machine learning technique where the goal is to predict an unordered set of elements. It is widely used in various applications such as object detection (Carion et al. 2020), point cloud classification (Qi et al. 2017), and multi-label classification (Yang et al. 2019). In the field of NER, Seq2Set (Tan et al. 2021) first reformulates the nested NER task as set prediction to eliminate order bias from the autoregressive model. Subsequently, PIQN (Shen et al. 2022) associates each entity with a learnable instance query to capture the location and type semantics. Recently, DQSetGen (Chen et al. 2024) extended set prediction to other sequence labeling tasks such as slot filling and part-of-speech (POS) Tagging. Different from these methods, we introduce the MQSPN, which exploits the set prediction paradigm in a new GMNER area to model the two-level relationship of multimodal entities.
Figure 2: (a). Overview of our MQSPN. MQS and MSP guide models to learn intra-entity and inter-entity relationships respectively. (b). The construction of Multi-grained Query Set (MQS) consists of Type-grained Queries and Learnable Entity-grained Queries (LEQ). (c). The detailed architecture of Query-guided Fusion Net (QFNet).
@@
Visual Grounding. Visual Grounding (VG) aims to detect the most relevant visual region based on a natural language query, i.e., phrase, category, or sentence. Most existing works can be divided into two branches. The first branch utilizes the one-stage detector, such as DETR (Carion et al. 2020) and YOLO (Redmon and Farhadi 2018), to directly output object bounding boxes in an end-to-end manner. The second branch first generates candidate regions with some region proposal methods, such as Region Proposal Network (RPN) (Girshick 2015) and selective search (Uijlings et al. 2013), and then selects the best-matching region based on language query. In this work, we follow the two-stage paradigm to construct an entity grounding model.
Our Method
Overview
As illustrated in Figure 2, we present a set prediction-based method named MQSPN with four different components, consisting of the Feature Extraction Module, Multi-grained Query Set (MQS), Query-guided Fusion Net (QFNet) and Multimodal Set Prediction Network (MSP). The objective of our MQSPN is to predict a set of multimodal quadruples which can be represented as:
$Y=\{\left(Y_{1}^{s},Y_{1}^{e},Y_{1}^{t},Y_{1}^{r}\right),\ldots,\left(Y_{m}^{s},Y_{m}^{e},Y_{m}^{t},Y_{m}^{r}\right)\},$ (1)
@@
where $(Y_{i}^{s},Y_{i}^{e},Y_{i}^{t},Y_{i}^{r})$ denote the $i$-th quadruple, $Y_{i}^{s}\in[0,n-1]$ and $Y_{i}^{e}\in[0,n-1]$ are the start and end boundary indices of the $i$-th target entity span. $Y_{i}^{t}$ refers to its corresponding entity type, and $Y_{i}^{r}$ denotes grounded region. Note that if the target entity cannot be grounded in the given image, $Y_{i}^{r}$ is $None$; otherwise, $Y_{i}^{r}$ consists of a 4-D spatial feature including the top-left and bottom-right coordinates of the grounded region.
Feature Extraction Module
@@
Text Representation. Given the input sentence $X$, textual encoder BERT (Devlin et al. 2019) is used to tokenize it into a sequence of word embeddings $H_{T}=([CLS],e_{1},...,e_{n},[SEP])$, where $e_{i}\in\mathbb{R}^{h},$ $h$ is the hidden dimension, $[CLS]$ and $[SEP]$ are special tokens of the beginning and end positions in word embeddings.
@@
Visual Representation. Given the input image $I$, we utilize VinVL (Zhang et al. 2021c) as the class-agnostic region proposal network (RPN) to obtain all candidate regions. Following the work of (Yu et al. 2023), we also retain the top-$k$ region proposals as our candidate regions, using the ViT-B/32 (Dosovitskiy et al. 2020) from pre-training CLIP (Radford et al. 2021) as vision encoder. The initial visual representation for the candidate regions are denoted as $\mathbf{V}=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\}$. To match those entities that are ungroundable in the images, we construct a special visual token embedding $\mathbf{v}_{[ug]}$ by feeding a blank image $\mathcal{T}$ into vision encoder. Finally, $\mathbf{v}_{[ug]}$ and $\mathbf{V}$ are concatenated to serve as the final visual representation $H_{V}\in\mathbb{R}^{(k+1)\times h}$, where $k$ is the number of candidate regions.
Multi-grained Query Set Construction
@@
Previous manually constructed query statements struggle to learn distinguishable features for different entities, hindering intra-entity connections modeling. In this section, we propose a learnable multi-grained query set to overcome this.
@@
$H_{Q}^{o}=BERT\left(Prompt\right)_{[\text{MASK}]}$ (2)
@@
Learnable Entity-grained Query Set. Entity-grained queries are randomly initialized as learnable input embeddings $H_{Q}^{e}\in\mathbb{R}^{u\times h}$. During training, the entity-grained semantics and corresponding relationships between candidate regions and entity spans can be learned automatically by these embeddings. To ensure that type-grained and entity-grained query embeddings have the same dimensions, we replicated the former $d$ times. Then the multi-grained query embedding is given by:
$H_{Q}\mathbf{}=H_{Q}^{e}\oplus\left[H_{Q}^{o}\right]^{d}$ (3)
@@
Query-guided Fusion Net
Previous multimodal fusion approaches suffered from direct fusion of textual and visual information due to semantic discrepancies and noisy information between multimodal data. Different from them, we use queries as intermediaries to guide the integration of textual representations and visual region representations respectively. This module includes three interaction mechanisms.
@@
Query-text Cross-attention Interaction. As shown in Figure 2 (b), the query set $H_{Q}$ and the textual sequence $H_{T}$ are fed into the transformer-based architecture (Vaswani et al. 2017). The cross-attention mechanism is used to fuse these unimodal features.
Query-region Prefix Integration. Following prefix tuning (Li and Liang 2021) and its successful applications (Chen et al. 2022a, b) in multimodal fusion, we propose query-region prefix integration to reduce the semantic bias between multimodal data. Details of the Query-region Prefix Integration are provided in the Appendix.